Reports 1-1 of 1 Clear search Modify search
CAL (General)
Shingo Hido - 19:12 Thursday 20 November 2025 (35612) Print this report
Calibration check after O4c
CAL group (11/19)
We performed calibration measurements to check the current parameters.

Estimated parameters in the previous measurements are as follows.
 H_etmxtm = 3.908232111e-14 @10Hz ( -0.34% from previous measurements)
 H_etmxim = 1.487776148e-14 @10Hz  ( -2.78% from previous measurements)
 Optical_gain = 1.891353135e+12   ( -2.48% from previous measurements)
 Cavity_pole = 18.376366917 Hz  ( -0.35% from previous measurements)

No significant deviations were observed compared to the previous measurements. We did not update the parameters.
Images attached to this report
Search Help
×

Warning

×